The attendance at a baseball game was 400 people. Student tickets cost $2 and adult tickets cost $3. Total ticket sales were $1050. How many tickets of each type were sold ?
Answer by addingup(3677)

The attendance at a baseball game was 400 people. Student tickets cost $2 and adult tickets cost $3. Total ticket sales were $1050. How many tickets of each
s+a= 400 therefore s= 400-a
2s+3a= 1050 Substituting:
400-a+3a= 1050
400+2a=1050
2a= 650
a= 325 is the $ in adult tickets
1050-325= 725 is the $ in student tickets